# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ Abstract
Cooperative and non-cooperative localization frequently arise together in wireless sensor networks, particularly when sensor positions are uncertain and targets are unable to communicate with the network. While joint processing can eliminate the delay in target estimation found in sequential approaches, it introduces complex variable coupling, posing challenges in both modeling and optimization. This paper presents a joint modeling approach that formulates cooperative and non-cooperative localization as a single optimization problem. To address the resulting coupling, we introduce auxiliary variables that enable structural decoupling and distributed computation. Building on this formulation, we develop the Scaled Proximal Alternating Direction Method of Multipliers for Joint Cooperative and Non-Cooperative Localization (SP-ADMM-JCNL). Leveraging the problem’s structured design, we provide theoretical guarantees that the algorithm generates a sequence converging globally to the Karush-Kuhn-Tucker (KKT) point of the reformulated problem and further to a critical point of the original non-convex objective function, with a sublinear rate of $\mathcal{O}(1/T)$. Experiments on both synthetic and benchmark datasets demonstrate that SP-ADMM-JCNL achieves accurate and reliable localization performance.
I Introduction
TABLE I: Comparison Between JCNL and SCNL Approaches
Aspect JCNL (Joint) SCNL (Separate)
Modeling Joint estimation of sensors and target Two-stage: estimation of sensors and then target
Optimization Single least-squares problem Sequential least-squares problems
@@ I Introduction
Information Flow Target data may aid sensor localization Used only after sensor localization
@@
Target Estimation Timeliness ✓ (estimated simultaneously with sensors) ✗ (target waits for sensor positions)
@@
Wireless Sensor Networks (WSNs), equipped with sensing, processing and communication capabilities, have been widely deployed in various domains such as military operations, environmental monitoring, healthcare, aerospace and industrial systems [1, 2, 3, 4, 5]. Localization serves as a fundamental enabler of the functionality of WSNs and typically involves two primary tasks: cooperative localization and non-cooperative localization. The former focuses on estimating the positions of sensor nodes using a limited number of anchors with precisely known locations and noisy range measurements between neighboring sensors [6]. The latter addresses the localization of external targets that are unable to communicate with the network, relying on passive observations collected by sensor nodes and their coordination with nearby nodes.
@@
Both cooperative and non-cooperative localization problems have been extensively studied in the context of wireless sensor networks (WSNs). These tasks typically rely on distance-related observations obtained through techniques such as time of arrival (TOA) [7, 8], time difference of arrival (TDOA) [9], and received signal strength (RSS) [10, 11]. Recently, optimization-based approaches have gained increasing attention due to their faster inference and theoretical guarantees on convergence. However, the corresponding maximum likelihood formulations usually result in non-convex and non-smooth least squares problems, which are difficult to solve efficiently and reliably [12]. To mitigate this, many existing methods adopt convex relaxation techniques that approximate the original problem and enable convergence to a global optimum of the relaxed version [13, 14, 15, 16, 17]. Nonetheless, such solutions may not correspond to critical points of the original non-convex formulation. In addition, most relaxation-based methods are implemented in centralized frameworks, where all measurements must be aggregated at a central node. This centralized architecture may increase communication overhead and introduce vulnerability to the malfunction of any single node.
@@
To overcome the limitations of centralized approaches, significant efforts have been devoted to developing distributed localization methods [18, 19]. Among them, the Alternating Direction Method of Multipliers (ADMM) has been recognized as a suitable tool for distributed convex optimization and has also shown promise in certain non-convex scenarios. In particular, ADMM has been proven extremely suitable for distributed convex problems due to its decomposability and convergence guarantees [20, 21, 22], and has also demonstrated effectiveness in tackling structured non-convex formulations under specific conditions. In the context of cooperative localization, a common strategy reformulates the original problem as the problem of two-block non-convex optimization with linear equality constraints. Based on this reformulation, distributed ADMM-based methods have been proposed in [23, 24] to solve the resulting problem. Building on these ideas, [25] introduced the Scaled Proximal ADMM (SP-ADMM) method—an efficient distributed algorithm with relatively low computational complexity—and established theoretical guarantees for its convergence and performance. In non-cooperative localization, [26] addresses a sensor network-based event localization problem using ADMM, under the assumption that the sensor positions are known.
@@
While cooperative and non-cooperative localization have each been extensively studied, they are typically handled separately. In many practical scenarios, both cooperative and non-cooperative localization problems arise simultaneously. This often occurs in settings such as battlefield or environmental monitoring, where sensor positions are uncertain and target cannot actively communicate with the network. A possible approach is to first perform cooperative localization and then estimate target positions based on the sensor estimates. For notational convenience, we refer to this sequential strategy as SCNL (Separate Cooperative and Non-Cooperative Localization). In contrast to this staged approach, joint approaches have shown clear advantages in various contexts. For instance, [27] investigates a joint solution to cooperative localization and multitarget tracking, while [28, 29, 30] explore the integration of cooperative localization with clock synchronization. Inspired by these, we propose a joint formulation— JCNL (Joint Cooperative and Non-Cooperative Localization)—that models and solves both localization tasks simultaneously. This integrated approach helps eliminate the target estimation delays inherent in sequential methods. Since the presence of a target often signals a potential risk or external threat, reducing such delays is particularly important in real-time applications.
@@
In this article, to retain the advantages of the SP-ADMM method in terms of communication efficiency and computational cost, and to address the modeling challenges introduced by the non-convex coupling between sensor and target variables, we develop a structured formulation of the joint localization problem. By introducing variable decoupling techniques, the original non-convex problem is transformed into a tractable form amenable to distributed optimization. Moreover, we establish the theoretical equivalence between the reformulated and original problems, ensuring that the solution accuracy is not compromised. In addition, our joint formulation naturally reduces to the standard non-cooperative localization problem when all sensors are treated as anchors, implying that the theoretical guarantees presented in this work also apply to that special case. A detailed comparison between JCNL and SCNL is provided in Table I.
@@
  • We propose a distributed least squares formulation that jointly addresses cooperative (sensor self-localization) and non-cooperative (target localization) tasks. Sensor and target positions are modeled together and estimated simultaneously under a unified optimization framework, instead of being decoupled into sequential stages. This joint formulation allows for consistent and concurrent estimation across heterogeneous node types, enhancing the coherence and accuracy of localization in distributed settings where sensor positions are uncertain and targets cannot actively communicate with the network.
  • We generalize the SP-ADMM algorithm to solve the more complex joint localization problem involving both sensors and passive targets. This extension is nontrivial, as the coupling between sensor and target variables introduces new structural challenges that prevent direct decomposition into simple update steps. To address this, we redesign several algorithmic components—including variable splitting and local update rules — so that the resulting distributed solver accommodates the jointly structured least squares model, maintains per-node computation tractability, and preserves scalability with respect to network size.
@@
TABLE II: Node Types in WSNs for Target Localization
Node Type Position Communication Enabled? Number
Sensor ($\mathcal{N}$) Anchor ($\mathcal{A}$) Known $m$
Agent ($\mathcal{N}\setminus\mathcal{A}$) Unknown NN
Target Unknown 11
@@
Section II presents the mathematical formulation of the joint localization problem. Section III introduces the proposed SP-ADMM-JCNL algorithm. Section IV provides the convergence analysis. Section V presents numerical evaluations, including comparisons with baseline methods. Finally, the conlusion is given in Section VI.
II Problem Formulation
II-A Notation
@@
where $\|\mathbf{z}\|_{\mathbf{W}}^{2}:=\langle\mathbf{z},\mathbf{Wz}\rangle$ denotes the $\mathbf{W}$-induced norm for $\mathbf{z}\in\mathbb{R}^{(4N_{i}+2)n}$. We also use $\delta_{\mathcal{C}}(\mathbf{x})$ to denote the indicator function of a closed convex set $\mathcal{C}$, defined as
@@
The set $\mathcal{C}$ typically represents a feasible region or constraint set imposed on the variables in the optimization problem.
II-B Problem Statement
We consider a sensor network for target localization, represented by an undirected connected graph $G=(\mathcal{N},\mathcal{E})$, where the network topology is assumed to be known. The sensor set is defined as $\mathcal{N}=\{1,2,\dots,N\}$, consisting of $N$ nodes. Among these, a subset of sensors serves as anchors with known true positions, collected in the set $\mathcal{A}=\{N-m+1,\dots,N\}\subset\mathcal{N}$. The remaining nodes are unknown-position sensors that require localization. The true position of node $i$ is denoted as $\mathbf{x}_{i}\in\mathbb{R}^{n}$, where $n$ represents the spatial dimension. For clarity, Table II summarizes the types of nodes in WSNs considered for target localization. We define the concatenated position vector as
$\mathbf{x}=[\mathbf{x}_{1}^{T},\mathbf{x}_{2}^{T},\dots,\mathbf{x}_{N}^{T}]^{T}.$
For each node $i\in\mathcal{N}$, we define its neighborhood set as
$\mathcal{N}_{i}=\{j\mid(i,j)\in\mathcal{E}\},$
where $N_{i}=|\mathcal{N}_{i}|$ represents its cardinality. If nodes $i$ and $j$ are neighbors, a communication link exists between them, enabling information exchange. Additionally, we assume that the true position of the target is given by $\mathbf{y}\in\mathbb{R}^{n}$. The noisy distance measurements between node $i$ and the target, as well as between node $i$ and its neighboring node $j\in\mathcal{N}_{i}$, are defined as follows:
$r_{i}$ $=\|\mathbf{x}_{i}-\mathbf{y}\|+\tau_{i},\quad i\in\mathcal{N},$ (1a)
@@ II-B Problem Statement
Here, $\tau_{i}$ and $\omega_{i,j}$ are zero-mean, independent Gaussian-distributed measurement noise terms. We assume that $d_{i,j}=d_{j,i}$.
The objective of this study is to estimate the positions of both the unknown sensors and the target based on the given measurement model. Using the notation defined above, the maximum likelihood estimator can be obtained by solving the following non-convex constrained optimization problem:
$\min_{\mathbf{x}\in\mathbb{R}^{nN},\mathbf{y}\in\mathbb{R}^{n}}$ $\quad\frac{1}{2}\sum_{i\in\mathcal{N}}f_{i}(\mathbf{x},\mathbf{y})$ (2a)
subject to $\quad\mathbf{x}_{k}=\mathbf{a}_{k},\quad\forall k\in\mathcal{A},$ (2b)
where
$f_{i}(\mathbf{x},\mathbf{y})=\sum_{j\in\mathcal{N}_{i}}(\|\mathbf{x}_{i}-\mathbf{x}_{j}\|-d_{i,j})^{2}+(\|\mathbf{x}_{i}-\mathbf{y}\|-r_{i})^{2},\forall i\in\mathcal{N}.$ (3)
[…]
II-C Problem Reformulation
Our initial step involves formulating an equivalent smooth and constrained reformulation of problem (2), which serves as the foundation for both the algorithm design and the convergence analysis in this study. It is important to note that the objective function in problem (2) can be explicitly expressed as

### v2

# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed SP-ADMM Algorithms

@@ Abstract
The integration of cooperative and non-cooperative localization is fundamentally important, as these two modes frequently coexist in wireless sensor networks—especially when sensor positions are uncertain and targets are unable to communicate with the network. This paper presents a joint modeling approach that formulates cooperative and non-cooperative localization as a single optimization problem. By processing both tasks jointly, the proposed method eliminates the latency inherent in sequential approaches that perform cooperative localization first, followed by non-cooperative localization. However, this joint formulation introduces complex variable coupling, posing challenges in both modeling and optimization. To address this coupling, we introduce auxiliary variables that enable structural decoupling and facilitate distributed computation. Building on this formulation, we develop the Scaled Proximal Alternating Direction Method of Multipliers for Joint Cooperative and Non-Cooperative Localization (SP-ADMM-JCNL). Leveraging the structured design of the problem, we provide theoretical guarantees that the algorithm generates a sequence converging globally to a Karush–Kuhn–Tucker (KKT) point of the reformulated problem, and further to a critical point of the original non-convex objective function, with a sublinear convergence rate of $O(1/T)$. Experiments demonstrate that SP-ADMM-JCNL achieves accurate and reliable localization performance.
keywords:
Joint localization; Consensus; Distributed algorithm; Scaled Proximal ADMM.
thanks: This paper was not presented at any IFAC meeting. Corresponding author Xiaojing Shen. e-mail: shenxj@scu.edu.cn
1 Introduction
Table 1: Comparison Between JCNL and SCNL Approaches
Aspect SCNL (Separate) JCNL (Joint)
Modeling Two-stage: sensors → targets Joint: simultaneous estimation
Optimization Sequential least-squares problems Single least-squares problem
@@ 1 Introduction
Information Flow Used only after sensor localization Targets data may aid sensor localization
@@
Targets Estimation Timeliness ✗ (targets wait for sensor positions) ✓ (estimated simultaneously with sensors)
@@
Wireless Sensor Networks (WSNs), equipped with sensing, processing and communication capabilities, have been widely deployed in various domains such as military operations, environmental monitoring, healthcare, aerospace and industrial systems [1, 2, 3, 4, 5]. Localization serves as a fundamental enabler of the functionality of WSNs and typically involves two primary tasks: cooperative localization and non-cooperative localization. The former focuses on estimating the positions of sensor sensors using a limited number of anchors with precisely known locations and noisy range measurements between neighboring sensors [6]. The latter addresses the localization of external targets that are unable to communicate with the network, relying on passive observations collected by sensor sensors and their coordination with nearby sensors.
@@
Both cooperative and non-cooperative localization problems have been extensively studied in the context of WSNs. These tasks typically rely on distance-related observations obtained through techniques such as time of arrival (TOA) [7, 8], time difference of arrival (TDOA) [9], and received signal strength (RSS) [10, 11]. Recently, optimization-based approaches have gained increasing attention due to their faster inference and theoretical guarantees on convergence. However, the corresponding maximum likelihood formulations usually result in non-convex and non-smooth least squares problems, which are difficult to solve efficiently and reliably [12]. To mitigate this, many existing methods adopt convex relaxation techniques that approximate the original problem and enable convergence to a global optimum of the relaxed version [13, 14, 15]. Nonetheless, such solutions may not correspond to critical points of the original non-convex formulation. In addition, most relaxation-based methods are implemented in centralized frameworks, where all measurements must be aggregated at a central sensor. This centralized architecture may increase communication overhead and introduce vulnerability to the malfunction of any single sensor.
@@
To overcome the limitations of centralized approaches, significant efforts have been devoted to developing distributed localization methods [16, 17]. Among them, the Alternating Direction Method of Multipliers (ADMM) has been recognized as a suitable tool for distributed convex optimization and has also shown promise in certain non-convex scenarios. In particular, ADMM has been proven extremely suitable for distributed convex problems due to its decomposability and convergence guarantees [18, 19, 20], and has also demonstrated effectiveness in tackling structured non-convex formulations under specific conditions. In the context of cooperative localization, a common strategy reformulates the original problem as the problem of two-block non-convex optimization with linear equality constraints. Based on this reformulation, distributed ADMM-based methods have been proposed in [21, 22] to solve the resulting problem. Building on these ideas, [23] introduced the Scaled Proximal ADMM (SP-ADMM) method—an efficient distributed algorithm with relatively low computational complexity—and established theoretical guarantees for its convergence and performance. More recently, graph neural network based approaches have also been explored for network localization [24, 25]. In non-cooperative localization, [26] addresses a sensor network-based event localization problem using ADMM, under the assumption that the sensor positions are known.
@@
While cooperative and non-cooperative localization have each been extensively studied, they are typically handled separately. In many practical scenarios, both cooperative and non-cooperative localization problems arise simultaneously. This often occurs in settings such as battlefield or environmental monitoring, where sensor positions are uncertain and targets cannot actively communicate with the network. A possible approach is to first perform cooperative localization and then estimate target positions based on the sensor estimates. For notational convenience, we refer to this sequential strategy as SCNL (Separate Cooperative and Non-Cooperative Localization). In contrast to this staged approach, joint approaches have shown clear advantages in various contexts. For instance, [27] investigates a joint solution to cooperative localization and multitarget tracking, while [28, 29, 30] explore the integration of cooperative localization with clock synchronization. Inspired by these, we propose a joint formulation— JCNL (Joint Cooperative and Non-Cooperative Localization)—that models and solves both localization tasks simultaneously. This integrated approach helps eliminate the target estimations delay inherent in sequential methods. Since the presence of targets often signals a potential risk or external threat, reducing such delays is particularly important in real-time applications.
@@
In this article, to retain the advantages of the SP-ADMM method in terms of communication efficiency and computational cost, and to address the modeling challenges introduced by the non-convex coupling between sensor and target variables, we develop a structured formulation of the joint localization problem. By introducing variable decoupling techniques, the original non-convex problem is transformed into a tractable form amenable to distributed optimization. Moreover, we establish the theoretical equivalence between the reformulated and original problems, ensuring that the solution accuracy is not compromised. In addition, our joint formulation naturally reduces to the standard non-cooperative localization problem when all sensors are treated as anchors, implying that the theoretical guarantees presented in this work also apply to that special case. A detailed comparison between JCNL and SCNL is provided in Table 1.
@@
  • We present a distributed least squares formulation that jointly addresses cooperative (sensor self-localization) and non-cooperative (targets localization) tasks. Sensor and passive target positions are modeled together and estimated simultaneously under a unified optimization framework, instead of being decoupled into sequential stages. This joint formulation allows for consistent and concurrent estimation across heterogeneous sensor types, enhancing the coherence and accuracy of localization in distributed settings where sensor positions are uncertain and targets cannot actively communicate with the network.
  • Inspired by the work in [23], we propose distributed SP-ADMM algorithms to address the more complex problem of joint localization involving both sensors and passive targets. The coupling between sensor and target variables introduces new structural challenges that prevent direct decomposition into simple update steps. To overcome this, we redesign several algorithmic components—including variable splitting and local update rules—so that the resulting distributed solver accommodates the jointly structured least squares model, maintains tractability for per-sensor computation, and preserves scalability with respect to network size.
@@
Table 2: Node Types in WSNs for Passive Targets Localization
Nodes Type Position Communication Enabled? Number
Sensor ($\mathcal{N}$) Anchor ($\mathcal{A}$) Known $\color[rgb]{0,0,0}A$}
Agent ($\mathcal{N}\setminus\mathcal{A}$) Unknown $\color[rgb]{0,0,0}N-A$}
Target($\mathcal{M}$) Unknown $\color[rgb]{0,0,0}M$}
@@
Section 2 presents the mathematical formulation of the joint localization problem. Section 3 introduces the proposed SP-ADMM-JCNL algorithm. Section 4 provides the convergence analysis. Section 5 presents numerical evaluations, including comparisons with baseline methods. Finally, the conlusion is given in Section 6.
2 Problem Formulation
2.1 Notation
@@
where $\|\mathbf{z}\|_{\mathbf{W}}^{2}:=\langle\mathbf{z},\mathbf{Wz}\rangle$ denotes the $\mathbf{W}$-induced norm for $\mathbf{z}$. We also use $\delta_{\mathcal{C}}(\mathbf{x})$ to denote the indicator function of a closed convex set $\mathcal{C}$, defined as
@@
The set $\mathcal{C}$ typically represents a feasible region or constraint set imposed on the variables in the optimization problem. Finally, we define with $blockdiag(\cdot)$ the block diagonal matrix whose diagonal contains the input blocks matrices.
2.2 Problem Statement
We consider a sensor network for localizing multiple passive targets, represented by an undirected connected graph $G=(\mathcal{N},\mathcal{E})$, where the network topology is assumed to be known. The sensor set $\mathcal{N}=\{1,2,\dots,N\}$ consists of $N$ sensors, some of which are anchors with known true positions collected in the set $\mathcal{A}=\{\mathbf{a}_{N-A+1},\dots,\mathbf{a}_{N}\}\subset\mathcal{N}$. The remaining sensors are unknown-position sensors that require localization. The true position of sensor $i$ is denoted as $\mathbf{x}_{i}\in\mathbb{R}^{n}$, where $n$ represents the spatial dimension. Additionally, we assume that there are $M$ passive targets, indexed by the set $\mathcal{M}=\{1,2,\dots,M\}$. The true position of target $k\in\mathcal{M}$ is denoted by $\mathbf{y}_{k}\in\mathbb{R}^{n}$. For clarity, Table 2 summarizes the types of sensors in WSNs considered for targets localization. We define the concatenated position vector as $\mathbf{x}=[\mathbf{x}_{1}^{T},\mathbf{x}_{2}^{T},\dots,\mathbf{x}_{N}^{T}]^{T},\mathbf{y}=[\mathbf{y}_{1}^{T},\mathbf{y}_{2}^{T},\dots,\mathbf{y}_{M}^{T}]^{T}.$ For each sensor $i\in\mathcal{N}$, we define its neighborhood set as $\mathcal{N}_{i}=\{j\mid(i,j)\in\mathcal{E}\}\subseteq\mathcal{N}$ with cardinality $N_{i}=|\mathcal{N}_{i}|$, where a communication link between neighboring sensors $i$ and $j$ enables information exchange; similarly, we define the set of targets within line-of-sight for sensor $i$ as $\mathcal{M}_{i}\subseteq\mathcal{M}$ with $M_{i}=|\mathcal{M}_{i}|$, allowing sensor $i$ to obtain noisy range measurements to any target ${\color[rgb]{0,0,0}k\in\mathcal{M}}_{i}$ by assuming that targets emit distinguishable signals for correct measurement association.
Figure 1: Exemplary illustration of the considered scenario.
In the considered scenario, two types of range measurements are available: sensor–sensor measurements obtained via communication links between neighboring sensors, and sensor–target measurements acquired through active sensing (e.g., radar or sonar) for targets within line-of-sight. Targets are passive and cannot perform ranging; only sensors can measure distances to targets. An exemplary illustration of this scenario is provided in Fig. 1. The noisy distance measurements between sensor $i$ and target $k\in\mathcal{M}_{i}$, as well as between sensor $i$ and its neighboring sensor $j\in\mathcal{N}_{i}$, are defined as follows:
$r_{i,{\color[rgb]{0,0,0}k}}$ $=\|\mathbf{x}_{i}-\mathbf{y}_{k}\|+\tau_{i,{\color[rgb]{0,0,0}k}},\quad i\in\mathcal{N},\;{\color[rgb]{0,0,0}k\in\mathcal{M}}_{i},$ (1a)
@@ 2.2 Problem Statement
Here, $\tau_{i,{\color[rgb]{0,0,0}k}}$ and $\omega_{i,j}$ are zero-mean, independent Gaussian-distributed measurement noise terms. We assume that $d_{i,j}=d_{j,i}$.
The objective of this study is to estimate the positions of both the unknown sensors and the multiple targets based on the given measurement model. Using the notation defined above, the maximum likelihood estimator can be obtained by solving the following non-convex constrained optimization problem:
$\min_{\mathbf{x}\in\mathbb{R}^{nN},\mathbf{y}\in\mathbb{R}^{nM}}$ $\quad\frac{1}{2}\sum_{i\in\mathcal{N}}f_{i}(\mathbf{x},\mathbf{y})$ (2a)
subject to $\quad\mathbf{x}_{l}=\mathbf{a}_{l},\quad\forall l\in\mathcal{A},$ (2b)
where $\forall i\in\mathcal{N},$
$f_{i}(\mathbf{x},\mathbf{y})=$ $\sum_{j\in\mathcal{N}_{i}}(\|\mathbf{x}_{i}-\mathbf{x}_{j}\|-d_{i,j})^{2}$
$+{\color[rgb]{0,0,0}\sum_{k\in\mathcal{M}_{i}}(\|\mathbf{x}_{i}-\mathbf{y}_{k}\|-r_{i,{\color[rgb]{0,0,0}k}})^{2}}.$ (3)
We note that when $\mathcal{M}=\emptyset$, the problem reduces to the classic cooperative sensor network localization problem; and when $\mathcal{A}=\mathcal{N}$, problem (2) becomes a pure non-cooperative localization task, where only the targets’ positions need to be estimated.
Remark 1.
Unlike conventional cooperative localization where all nodes are cooperative, our framework distinguishes between active sensors (cooperative) and passive targets (non-cooperative). While sensors refine their own positions through inter-node communication and range updates, targets lack both the sensing and communication capabilities to participate in the network. Consequently, a target’s position is not self-determined but is jointly estimated by its observing sensors. This necessitates a decoupled formulation where target localization relies on consensus constraints across multiple sensors (see Section 2.3 for detailed discussion). Beyond the standard spatial coupling via pairwise measurements $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|$, any sensors $i$ and $j$ observing a common target $k\in(\mathcal{M}_{i}\cap\mathcal{M}_{j})$ become implicitly coupled through the shared target variable $\mathbf{y}_{k}$. This multi-path architecture enables sensors to pool mutual geometric information, yet it intensifies the coupling relationships, posing challenges for distributed parallel processing.
2.3 Problem Reformulation
Our initial step involves formulating an equivalent smooth and constrained reformulation of problem (2), which serves as the foundation for both the algorithm design and the convergence analysis in this study. Note that the objective function in (2) can be explicitly rewritten for all $i\in\mathcal{N}$ as follows: